OpenStudy (anonymous):

Solve for x. 3(3x - 1) + 2(3 - x) = 0 a. x = 3/7 b. x = -3/7 c. x = 1/3 d. x = -1/3

OpenStudy (anonymous):

Simplifying 3(3x + -1) + 2(3 + -1x) = 0 Reorder the terms: 3(-1 + 3x) + 2(3 + -1x) = 0 (-1 * 3 + 3x * 3) + 2(3 + -1x) = 0 (-3 + 9x) + 2(3 + -1x) = 0 -3 + 9x + (3 * 2 + -1x * 2) = 0 -3 + 9x + (6 + -2x) = 0 Reorder the terms: -3 + 6 + 9x + -2x = 0 Combine like terms: -3 + 6 = 3 3 + 9x + -2x = 0 Combine like terms: 9x + -2x = 7x 3 + 7x = 0 Solving 3 + 7x = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3' to each side of the equation. 3 + -3 + 7x = 0 + -3 Combine like terms: 3 + -3 = 0 0 + 7x = 0 + -3 7x = 0 + -3 Combine like terms: 0 + -3 = -3 7x = -3 Divide each side by '7'. x = -0.4285714286 Simplifying x = -0.4285714286 Which is B(: